a pilot flew to a parcel of land and back in 5 ho
urs. The rate out was 150 mph and the rate returning was 100 mph how far was the parcel of land

Answer by Alan3354(69443)

a pilot flew to a parcel of land and back in 5 ho
urs. The rate out was 150 mph and the rate returning was 100 mph how far was the parcel of land
-----------
Avg speed of the round-trip = 2*150*100/(150+100) = 120 mi/hr
RT distance = 5*120 = 600 miles
--> 300 miles each way